# Rates of Convergence in the Central Limit Theorem for Markov Chains, with an Application to TD Learning

### v2

@@ 3 Markov Chain Central Limit Theorem
It is worth noting that, if we set $\beta=1-2/\log n,$ we require third moments to exist for the non-asymptotic martingale central limit theorem to hold. However, this condition is easily satisfied if we assume the Markov chain has a finite state space and is irreducible and aperiodic, as we will do so in the next section. If one were to apply Theorems 1 and 2 to more general scenarios, then the third moment condition must be verified or assumed.
@@ 4 An Application to TD Learning
Temporal Difference (TD) Learning is a common method to learn the performance of a policy in reinforcement learning; see [50, 1] for details. In [2], it was shown that TD learning can be represented as a jump Markovian linear dynamical system as follows:
@@
    $\{X_{k}\}$ is an irreducible, aperioidic, finite state Markov chain. So it satisfies Assumption 2, item (1).
@@
Our main result of this section characterizes the rate of convergence of this algorithm for an appropriately chosen $\epsilon$. As in the previous theorem, we have chosen to write the result using the $\mathcal{O}$ notation rather than making the constants explicit, but these constants can be made explicit by following the details the proof of the theorem.
@@
$d_{\mathcal{W}}(\sqrt{n}(\bar{\theta}_{n}-\theta^{*}),(\bar{A}^{-1}\Sigma_{\infty}\bar{A}^{-T})^{1/2}Z)=\mathcal{O}\Bigl{(}\max\left(\frac{\sqrt{\log n}}{n^{\delta-0.5}},\frac{1}{n^{1-\delta}}\right)\Bigr{)},$
@@
Properties of $\Phi_{j}^{n}$ are discussed in [14] and $\Upsilon_{j}^{n}=\frac{\epsilon_{j}}{\epsilon_{j+1}}\Phi_{j}^{n+1}.$ Then, we have
@@
and, from [14], $\Phi_{0}^{n}$ (and hence, $\Upsilon_{0}^{n}$) is bounded independent of $n$, we conclude that this term is $\mathcal{O}(\frac{1}{\sqrt{n}}).$
@@
where in the last step we have used the fact that solution to the matrix Poisson solution is bounded due to the assumptions on our Markov chain and $E(||\Delta_{n}||^{2})$ is bounded from [4]. Next, due to the fact that our Markov chain’s state space is finite, there exists constants $\kappa_{3}$ and $\kappa_{4}$ such that
@@
for some constant $\kappa_{6}.$ It is known from [14] that this term goes to zero as $n\rightarrow\infty$. In Appendix B, we show that this term is $\mathcal{O}(1/n^{1-\delta}).$
@@
Let $\Xi_{i},$ $i=1,2,3,4$ be the terms in (10)-(13) and $\Xi_{5}$ be (14). The different steps of the proof have shown that each of $E(||\Xi_{i}||)$ for $i=1,2,3,4$ is either $\mathcal{O}(\frac{\sqrt{\log n}}{n^{\delta-0.5}})$ or $\mathcal{O}(\frac{1}{n^{1-\delta}})$. Now we apply Theorem 2 to $\Xi_{5}.$ Since $\Xi_{5}$ converges to a Gaussian distribution in the Wasserstein distance at rate $\mathcal{O}(1/n^{\beta})$ and $\beta<1$ can be arbitrary due to our assumptions on the Markov chain, we have proved the theorem.
@@ 5 Conclusions
  • One can try to use the martingale CLT results in this paper to strengthen the results and allow for Markovian noise in [24]. However, these extensions would require us to go beyond the linear setting in this paper and consider nonlinear problems. Along similar lines, it would be also interesting to see if our results can be used to study contractive stochastic approximation as in [43] or other stochastic approximation schemes [51, 52].
  • Under some conditions, the rate of convergence in the central limit theorem for i.i.d. random vectors is $\mathcal{O}(1/\sqrt{n})$ [26]. It would be interesting to see if the $\mathcal{O}(\log n/\sqrt{n})$ rate of convergence in Corollary (1) can be improved to $\mathcal{O}(1/\sqrt{n})$ under possibly additional conditions than the ones used in this paper. The Edgeworth expansion for scalar functions of Markov chains in [53] suggests that the $\log n$ could possibly be eliminated for Markov chains as well. The reason for $\log n$ in our result is due to the fact that we use the regularity results for the solution to Stein’s equation in [25] after we do the Lindeberg decomposition in the proof of the rate of convergence of martingale CLT. A different approach may be required to eliminate the $\mathcal{O}(\log n)$ term in the numerator.
  • The Wassertein distance used in this paper is the $1$-Wasserstein distance, it would be interesting to obtain convergence rates in the stronger $p$-Wasserstein distance. Work along these lines under different conditions than in this paper can be found in [54], it would be interesting to see if such results can be obtained for martingales and Markov chains.
@@
  • Ignoring the $\log$ term in Theorem‘3, our result suggests that $0.75$ is a good choice for $\delta$. This is consistent with the conclusion in [20]. However, this ignores the constants hidden in the $\mathcal{O}$ which could depends on $\delta$ (see Appendix B). Moreover, what we have is only an upper bound, so a more refined analysis is needed to understand a good choice for $\delta.$
@@ Acknowledgment:
  • I thank Prof. Siva Theja Maguluri, Georgia Tech, for many stimulating discussions and drawing my attention to several references during the course of this work, Prof. Sean Meyn, University of Florida, for pointing out the bias issue with fixed stepsize TD learning and why it cannot be eliminated by averaging, and Prof. Mehrdad Moharrami, University of Iowa, for carefully reading an earlier version of the paper and providing many useful comments.
@@
In both this section and the next, notation that we will use for the constants are just for the particular section of the appendix alone and so we may occasionally reuse some letters that were used for other constants elsewhere in the paper for constants here. We start with the following result from [4, Theorem 12]: there exists a positive integer $k^{*}$ such that for any $k\geq\hat{k},$ we have
@@ Appendix B Properties of $\Upsilon_{j}^{n}$ and $\Phi_{j}^{n}$
Recall the definition of $\Upsilon_{j}^{n}$ and $\Phi_{j}^{n}$ in (9). In this section of the appendix, we examine certain expressions which are shown be either bounded or converging to zero in [14], and derive more explicit bounds for them as a function of $\delta$.
From [14, page 846], we know that $\Phi_{j}^{t}$ can be written as

### v3

@@ 3.1 Finite State Space Markov Chains
As in (Gallouët et al. 2018), we can obtain the following corollary by choosing $\beta=1-2/\log n$ and noting that $\tilde{C}_{1}$ and $\tilde{C}_{2}$ become $O(\log n)$ in that case. It is worth noting that if we set $\beta=1-2/\log n,$ we require third moments to exist for the non-asymptotic martingale central limit theorem to hold. This condition is satisfied since the solution to Poisson’s equation is bounded under the assumption that the Markov chain has a finite state space and is irreducible and aperiodic. $\Box$
3.2 Extension to General State-Space Markov Chains
We now extend the results to general state-space Markov chains. The ideas are similar to those of the finite state-space case, but we need additional assumptions and work to prove the result. To avoid unnecessary repetition, we use much of the same notation as in the previous section, while additional terminology and machinery are borrowed from (Meyn and Tweedie 2012). Not surprisingly, we need stronger assumptions than the conditions needed in (Meyn and Tweedie 2012) to establish the CLT since we are interested in obtaining the rate of convergence.
Theorem 3.
Let $X$ be a $\psi$-irreducible, aperiodic Markov chain over a state space $\mathcal{S}$. Suppose there exists a Lyapunov function $\mathscr{V}\geq 1$, with $E(V(X_{0}))<\infty$, such that the following drift condition is satisfied:
$E(\mathscr{V}(X_{k+1})|X_{k})\leq\lambda\mathscr{V}(X_{k})+L$ (9)
for some $\lambda\in[0,1)$ and $L\in[0,\infty).$ (Note that it is well known that one can equivalently assume a traditional Foster-type drift condition and strengthen it to (9) by imposing reasonable conditions on the increments of the Lyapunov function; see (Popov 1977, Hajek 1982, Spieksma and Tweedie 1994) and (Meyn and Tweedie 2012, Theorem 16.3.1).) Then, the following results hold:
  1. 1.
    For any $\beta\in(0,1)$, if $|r(x)|^{2+\beta}\leq\mathscr{V}(x)\,\forall x\in S,$ then
    $d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{1}{(1-\beta)n^{\beta/2}}\Bigr{)},$
    where $Z\sim N(0,I).$
  2. 2.
    If $|r(x)|^{3}\leq\mathscr{V}(x)\,\forall x\in S,$ then
    $d_{\mathcal{W}}(U_{n},\Sigma_{\infty}^{1/2}Z)=\mathcal{O}\Bigl{(}\frac{\log n}{\sqrt{n}}\Bigr{)}.$
Proof.
We will only present the additional steps needed beyond the finite-state case to obtain the second statement in the theorem. The proof of the first statement is nearly identical.
Noting the decomposition in (7) and as in the rest of the proof of Theorem 2, we need to prove the following four facts:
  1. 1.
    $E(||m_{k}||^{3}$) is uniformly bounded.
  2. 2.
    $E(||V(X_{0})-V(X_{n})||)$ is bounded uniformly in $n.$
  3. 3.
    $\Sigma_{\infty}=E_{\pi}((V(X_{1})-E(V(X_{1})|X_{0}))(V(X_{1})-E(V(X_{1})|X_{0}))^{T}$ exists, where $E_{\pi}$ denotes that $X_{0}\sim\pi.$
  4. 4.
    $Tr\Bigl{(}A_{k}\Bigl{(}\Sigma_{\infty}^{-1/2}E(\Sigma_{k})\Sigma_{\infty}^{-1/2}-I\Bigr{)}\Bigr{)}$ decays sufficiently fast as a function of $k.$
Once these facts are proved, the proof of the theorem is identical to the proof of Theorem 2.
Proof of (1):
From (9), we have:
$E(\mathscr{V}^{1/3}(X_{k+1})|X_{k})$ $\leq[E(\mathscr{V}(X_{k+1})|X_{k})]^{1/3}$
$\leq(\lambda\mathscr{V}(X_{k})+L)^{1/3}$
$\leq\lambda^{1/3}\mathscr{V}^{1/3}(X_{k})+L^{1/3},$
where the first inequality is Jensen’s inequality, the second is the from the drift condition (9) and the third can be seen by cubing the expressions in the second and third lines. Equivalently,
$E(\mathscr{V}^{1/3}(X_{k+1})|X_{k})-\mathscr{V}^{1/3}(X_{k})\leq-(1-\lambda^{1/3})\mathscr{V}^{1/3}(X_{k})+L^{1/3}.$ (10)
Now using the bound on $|r_{i}(x)|$ and Theorem 17.4.2 (and the discussion prior to it) in (Meyn and Tweedie 2012), we can conclude that there exists a solution to Poisson’s equation which satisfies the following bound:
$|V_{i}(x)|\leq B(\mathscr{V}^{1/3}(x)+1)\Rightarrow|V_{i}(x)|^{3}\leq 8B^{3}(\mathscr{V}(x)+1)\quad\forall x,$
where $V_{i}$ is the $i^{\rm th}$ component of $V$ and we have used the fact $(a+b)^{3}\leq(2\max(a,b))^{3}\leq 8(a^{3}+b^{3})$ for positive $a$ and $b.$
Next, considering the $i^{\rm th}$ component of $m_{k},$ we have
$|m_{ki}|^{3}$ $\leq 8(|E(V_{i}(X_{k}))|^{3}+|E(V_{i}(X_{k})|X_{k-1})|^{3}$
$=8(E(V_{i}^{3}(X_{k}))+E(V_{i}^{3}(X_{k})|X_{k-1}).$
Letting $E(\mathscr{V}(X_{\infty}))$ denote the expectation of $\mathscr{V}$ with respect to the stationary distribution, we have
$E(|m_{ki}|^{3})\leq 16E(V^{3}(X_{k}))$ $\leq 144B^{3}(E(\mathscr{V}(X_{k}))+1)$
$\leq 144B^{3}\left(E(\mathscr{V}(X_{\infty}))+1+|E(\mathscr{V}(X_{k}))-E(\mathscr{V}(X_{\infty}))|\right)$
$\leq 144B^{3}\left(E(\mathscr{V}(X_{\infty}))+1+RE(V(X_{0}))\rho^{k}\right),$
for some $\rho\in[0,1),$ $R\in[0,\infty)$, where the last step follows from the geometric ergodicity theorem (Meyn and Tweedie 2012, Theorem 15.0.1). Furthermore, from the $f$-norm ergodic theorem (Meyn and Tweedie 2012, Theorem 14.0.1), we have $E(V(X_{\infty}))<\infty,$ which proves that $E(||m_{k}||^{3})$ is uniformly bounded in $k.$
Proof of (2):
It suffices to show that $E(||V(X_{n})||)$ is uniformly bounded.
$E(|V_{i}(X_{n})|)\leq BE(\mathscr{V}^{1/3}(X_{n})+1)$ $\leq BE(\mathscr{V}(X_{n})+1)$
$\leq B\Big{(}|E(\mathscr{V}(X_{n})-\mathscr{V}(X_{\infty}))|+|E(\mathscr{V}(X_{\infty}))|+1\Big{)}.$
As discussed above, $|E(\mathscr{V}(X_{n})-\mathscr{V}(X_{\infty}))|$ is uniformly bounded in view of the geometric ergodic theorem, thus concluding the proof.
Proof of (3):
Consider the $(i,j)^{\rm th}$ element of $\Sigma_{\infty}:$
$Cov_{\pi}(E(V_{i}(X_{1}|X_{0}),E(V_{j}(X_{1}|X_{0}))\leq\sqrt{Var_{\pi}(E(V_{i}(X_{1})|X_{0}))Var_{\pi}(E(V_{i}(X_{1})|X_{0}))}.$
Now
$Var_{\pi}(E(V_{i}(X_{1})|X_{0})\leq E_{\pi}(E(V_{i}(X_{1})|X_{0}))^{2}$ $\leq B^{2}E_{\pi}(E(\mathscr{V}^{1/3}(X_{1})|X_{0})+1)^{2}$
$\leq 2B^{2}(E_{\pi}((E(\mathscr{V}^{1/3}(X_{1})|X_{0}))^{2}+1)$
$\leq 2B^{2}(E_{\pi}((E(\mathscr{V}(X_{1})|X_{0}))^{2/3})+1)$
$\leq 2B^{2}(E_{\pi}(E(\mathscr{V}(X_{1})|X_{0}))+1)$
$\leq 2B^{2}(E_{\pi}(\mathscr{V}(X_{1}))+1),$
which is finite, thus proving (3).
Proof of (4):
Consider the absolute value of the expectation of the $(i,j)^{\rm th}$ term of $\Sigma_{k}=E(m_{k}m_{k}^{T}|X_{k-1}):$
$|Cov(E(V_{i}(X_{k})|X_{k-1}),E(V_{j}(X_{k})|X_{k-1})|X_{k})|$ $\leq B^{2}E\Big{(}(E(\mathscr{V}^{1/3}(X_{k})|X_{k-1})+E(\mathscr{V}^{1/3}(X_{k})))^{2}\Big{)}$
$\leq 2B^{2}E\Big{(}(E(\mathscr{V}^{2/3}(X_{k})|X_{k-1})+E(\mathscr{V}^{2/3}(X_{k})))\Big{)}$
$\leq 2B^{2}E\Big{(}E(\mathscr{V}(X_{k})|X_{k-1})+E(\mathscr{V}(X_{k}))\Big{)}$
$=4B^{2}E(\mathscr{V}(X_{k}))$
$=4B^{2}\Big{(}|E(\mathscr{V}(X_{k}))-E(\mathscr{V}(X_{\infty}))|+|E(\mathscr{V}(X_{\infty}))|\Big{)}$
Thus, the absolute value of the $(i,j)^{\rm th}$ term of $\Sigma_{k}-\Sigma_{\infty}$ is upper bounded by a constant by geometric ergodicity. $\Box$
@@ 4 An Application to TD Learning
Temporal Difference (TD) Learning is a common method to learn the performance of a policy in reinforcement learning; see (Bertsekas and Tsitsiklis 1996, Sutton and Barto 2018) for details. In (Tsitsiklis and Van Roy 1996), it was shown that TD learning can be represented as a jump Markovian linear dynamical system as follows:
@@
    $\{X_{k}\}$ is an irreducible, aperiodic, finite state Markov chain. So it satisfies Assumption 2, item (1).
@@
The first item in the assumption is standard in the reinforcement learning literature (Sutton and Barto 2018), while the second item has been established for TD learning in (Tsitsiklis and Van Roy 1996).
@@
Our main result of this section characterizes the rate of convergence of this algorithm for an appropriately chosen $\{\epsilon_{k}\}$. As in the previous theorem, we have chosen to write the result using the $\mathcal{O}$ notation rather than making the constants explicit, but these constants can be made explicit by following the details the proof of the theorem.
@@
$d_{\mathcal{W}}(\sqrt{n}(\bar{\theta}_{n}-\theta^{*}),(\bar{A}^{-1}\Sigma_{\infty}\bar{A}^{-T})^{1/2}Z)=\mathcal{O}\Bigl{(}\max\left(\frac{\sqrt{\log n}}{n^{\delta-0.5}},\frac{1}{n^{(1-\delta)/2}}\right)\Bigr{)},$
@@
Properties of $\Phi_{j}^{n}$ are discussed in (Polyak and Juditsky 1992) and $\Upsilon_{j}^{n}=\frac{\epsilon_{j}}{\epsilon_{j+1}}\Phi_{j}^{n+1}.$ Then, we have
@@
and, from (Polyak and Juditsky 1992), $\Phi_{0}^{n}$ (and hence, $\Upsilon_{0}^{n}$) is bounded independent of $n$, we conclude that this term is $\mathcal{O}(\frac{1}{\sqrt{n}}).$
@@
where in the last step we have used the fact that solution to the matrix Poisson solution is bounded due to the assumptions on our Markov chain and $E(||\Delta_{n}||^{2})$ is bounded from (Srikant and Ying 2019). Next, due to the fact that our Markov chain’s state space is finite, there exists constants $\kappa_{3}$ and $\kappa_{4}$ such that
@@
for some constant $\kappa_{6}.$ It is known from (Polyak and Juditsky 1992) that this term goes to zero as $n\rightarrow\infty$. In Appendix B, we show that this term is $\mathcal{O}(1/n^{1-\delta}).$
@@
Let $\Xi_{i},$ $i=1,2,3,4$ be the terms in (13)-(16) and $\Xi_{5}$ be (17). The different steps of the proof have shown that each of $E(||\Xi_{i}||)$ for $i=1,2,3,4$ is either $\mathcal{O}(\frac{\sqrt{\log n}}{n^{(\delta-0.5})/2})$ or $\mathcal{O}(\frac{1}{n^{1-\delta}})$. Now we apply Theorem 2 to $\Xi_{5}.$ Since $\Xi_{5}$ converges to a Gaussian distribution in the Wasserstein distance at rate $\mathcal{O}(1/n^{\beta})$ and $\beta<1$ can be arbitrary due to our assumptions on the Markov chain, we have proved the theorem. $\Box$
@@ 5 Conclusions
  • One can try to use the martingale CLT results in this paper to strengthen the results and allow for Markovian noise in (Anastasiou et al. 2019). However, these extensions would require us to go beyond the linear setting in this paper and consider nonlinear problems. Along similar lines, it would be also interesting to see if our results can be used to study contractive stochastic approximation as in (Chen et al. 2010) or other stochastic approximation schemes (Borkar 2009, Kushner and Yin 2003).
  • Under some conditions, the rate of convergence in the central limit theorem for i.i.d. random vectors is $\mathcal{O}(1/\sqrt{n})$ (Fang et al. 2019). It would be interesting to see if the $\mathcal{O}(\log n/\sqrt{n})$ rate of convergence in Theorem 1 can be improved to $\mathcal{O}(1/\sqrt{n})$ under possibly additional conditions than the ones used in this paper. The Edgeworth expansion for scalar functions of Markov chains in (Kontoyiannis and Meyn 2003) suggests that the $\log n$ could possibly be eliminated for Markov chains as well. The reason for $\log n$ in our result is due to the fact that we use the regularity results for the solution to Stein’s equation in (Gallouët et al. 2018) after we do the Lindeberg decomposition in the proof of the rate of convergence of martingale CLT. A different approach may be required to eliminate the $\mathcal{O}(\log n)$ term in the numerator.
  • The approach in (Gallouët et al. 2018) that we have used is known to be not tight in terms of the dependence on the dimension $d$ for i.i.d. random vectors. Under some conditions, the dependence on $d$ can be considerably strengthened in the i.i.d. case (Courtade et al. 2019, Zhai 2018). Whether that is possible for martingales and Markov chains is an interesting question for further research.
  • The Wassertein distance used in this paper is the $1$-Wasserstein distance, it would be interesting to obtain convergence rates in the stronger $p$-Wasserstein distance. Work along these lines under different conditions than in this paper can be found in (Bonis 2020), and in the previously mentioned references (Courtade et al. 2019, Zhai 2018). It would be interesting to see if such results can be obtained for martingales and Markov chains.
@@
  • Ignoring the $\log$ term in Theorem 4, our result suggests that $2/3$ is a good choice for $\delta$. This is different than the conclusion in (Haque et al. 2023), where their upper bound implies $\delta=0.75$ is the best choice. However, these bounds ignores the constants hidden in the $\mathcal{O}$ which could depends on $\delta$ (see Appendix B). Moreover, what we have is only an upper bound, so a more refined analysis is needed to understand a good choice for $\delta.$
@@ Acknowledgment:
  • I thank Prof. Siva Theja Maguluri, Georgia Tech, for many stimulating discussions and drawing my attention to several references during the course of this work, Prof. Sean Meyn, University of Florida, for pointing out the bias issue with fixed stepsize TD learning and why it cannot be eliminated by averaging, and other discussions, Prof. Mehrdad Moharrami, University of Iowa, for carefully reading an earlier version of the paper and providing many useful comments, and Weichen Wu, CMU, for pointing out a missing factor in Theorem 4 in an earlier version of the paper.
@@
In both this section and the next, notation that we will use for the constants are just for the particular section of the appendix alone and so we may occasionally reuse some letters that were used for other constants elsewhere in the paper for constants here. We start with the following result from (Srikant and Ying 2019, Theorem 12): there exists a positive integer $k^{*}$ such that for any $k\geq\hat{k},$ we have
@@ Appendix B Properties of $\Upsilon_{j}^{n}$ and $\Phi_{j}^{n}$
Recall the definition of $\Upsilon_{j}^{n}$ and $\Phi_{j}^{n}$ in (12). In this section of the appendix, we examine certain expressions which are shown be either bounded or converging to zero in (Polyak and Juditsky 1992), and derive more explicit bounds for them as a function of $\delta$.
From (Polyak and Juditsky 1992, page 846), we know that $\Phi_{j}^{t}$ can be written as